OpenStudy (anonymous):

The equaqtion \(ax^3+bx^2+cx+d\) is known to’ve three distinct real roots .How many roots are there of the equation : \(4(ax^3+bx^2+cx+d)(3ax+b)=(3ax^2+2bx+c)^2\)

OpenStudy (vishweshshrimali5):

Hi! Back again with a good question ?

OpenStudy (anonymous):

:)

OpenStudy (vishweshshrimali5):

How do you think we should solve this question?

OpenStudy (vishweshshrimali5):

Any idea ?

OpenStudy (vishweshshrimali5):

Okay one thing that I have in my mind is this: If the roots of the polynomial are say \(x_1, ~x_2, ~x_3\). Then, \[\large{ax^3 + bx^2 + cx + d = a(x-x_1)(x-x_2)(x-x_3)}\]

OpenStudy (vishweshshrimali5):

You can also use calculus to solve this problem but I would discuss it later on

OpenStudy (vishweshshrimali5):

Now plugging in this in the problem we can come up with something

OpenStudy (vishweshshrimali5):

Also, we know that: \[\large{\sum x_1 = -\cfrac{b}{a}}\]

OpenStudy (vishweshshrimali5):

\[\large{\sum x_1 x_2 = \cfrac{c}{a}}\] \[\large{x_1x_2x_3 = -\cfrac{d}{a}}\]

OpenStudy (vishweshshrimali5):

Now lets plug in the polynomial in the equation given in the problem

OpenStudy (vishweshshrimali5):

Or better would be this: \[\large{b = -a\sum x_1} \tag{1}\] \[\large{c = a\sum x_1 x_2}\tag{2}\] \[\large{d = -ax_1x_2x_3}\tag{3}\]

OpenStudy (vishweshshrimali5):

\[\Large 4(ax^3+bx^2+cx+d)(3ax+b)=(3ax^2+2bx+c)^2\] \[\large{4(3a^2x^4+4abx^3+x^2(b^2+3ac)+x(bc+3ad)+bd) = (3ax^2+2bx+c)^2}\]

OpenStudy (vishweshshrimali5):

\[{4(3a^2x^4+4abx^3+x^2(b^2+3ac)+x(bc+3ad)+bd) = (3ax^2+2bx+c)^2}\]

OpenStudy (vishweshshrimali5):

Simplifying this using wolfram

OpenStudy (vishweshshrimali5):

\[\large{6a^2x^4+8abx^3+x^2(3ac+3b^2)+x(3bc-3ad)+c^2-bd= 0}\]

OpenStudy (vishweshshrimali5):

*sigh* I am tired :D

OpenStudy (vishweshshrimali5):

Lets plug in the values of b, c and d from (1), (2) and (3)

OpenStudy (vishweshshrimali5):

Use the De' Carte's condition of real (positive and negative roots) We know that if there are 2n sign changes in f(x) starting from the first term, then it may have 2n or 2n-2 or 2n-4 so on upto 0 number of POSITIVE roots.And if there are 2n+1 sign changes then there could be 2n+1 or 2n-1 or 2n-3 so on up to 1 positive roots. And applying the same thing to f(-x) gives the number of negative roots.

OpenStudy (vishweshshrimali5):

I am just trying to come up with something :)

OpenStudy (vishweshshrimali5):

I am thinking of using the De' Carte's condition for ax^3 + bx^2 + cx + d first and come up with the signs of a,b,c and d

OpenStudy (vishweshshrimali5):

Any ideas anyone ?

OpenStudy (vishweshshrimali5):

@hartnn any idea ?

hartnn (hartnn):

seems like a B^2 -4AC =0 format (ax^3+bx^2+cx+d) X^2 + (3ax^2+2bx+c)X + (3ax+b) =0 has EQUAL roots....

OpenStudy (vishweshshrimali5):

Sure? I have not checked actually

OpenStudy (vishweshshrimali5):

Yeah correct..

OpenStudy (vishweshshrimali5):

Not sure but I think the answer may be 2 solutions. @JungHyunRan can you check it?

OpenStudy (vishweshshrimali5):

Did it work @JungHyunRan ?

OpenStudy (vishweshshrimali5):

I think (and am pretty sure) that the number of roots = 2 (just an intution) Also, I think the problem is not correctly stated or something is missing.

OpenStudy (anonymous):

yeah! It has 2 solutions Good job! @vishweshshrimali5 :)

OpenStudy (vishweshshrimali5):

Thanks :) It worked !! :D

OpenStudy (vishweshshrimali5):

Another of your tricky questions @JungHyunRan . I really appreciate your trying to solve questions which increase the skills of thinking out of the box which is the most required skill for betterment in maths. Best of luck!

OpenStudy (anonymous):

^_^ thank you